• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jmaxwell

    Log entry time 21:27:40 on April 13, 2012

    Entry number 374068

    keyword=Run Status, Plans

    We are continuing with production data, although today has been a slog.  
    A beam trip due to a fuse on the chicane magnet's safety lamp cost about 
    half an hour, but returning good, centered beam after this trip was 
    problematic.  In all these problems cost about 3 hours.
    
    Production data continued in spite of high rate in the 3rd arm.  This was 
    likely due to beam scraping the beam pipe somewhere, as the plots of the 
    slow raster at the target were clean.  
    
    Jixie and Chao worked for a couple hours after 18:00 to adjust the slow 
    raster size and beam position to avoid the beam pipe and keep the beam 
    off the target cup walls.  Fortunately, this work was done while taking 
    production data, so it caused minimal delay.  We are now running with the 
    beam offset to the right by 2mm to mostly avoid the pipe, while keeping 
    the raster open to a total 2cm to ensure the best possible polarization.
    
    We will continue with production through the night.  We should keep a 
    close eye on the target polarization, and scale the number of events for 
    each point to follow the polarization, as described in the run plan.
    
    The target polarization is still above 30%, so we will plan for an anneal 
    early tomorrow morning.  We will perform a tungsten calorimeter 
    calibration at the same time, and if there is time after that before the 
    anneal is done, we will have a quick access to allow a change to the 
    parasitic detectors.
    
    Moving forward through next week, we will polarize the target only 
    positively to take advantage of the higher polarization that this 
    configuration offers for this time-strapped energy setting.
    
    Day Summary:
    0:00-10:30: Production
    10:30-11:00: Access to fix chicane trip
    11:00-14:00: Restore beam, centering
    14:00-now: Production with slow raster and beam position study